a weak acid ha,having a ka of 1 x 10-9,ionises in water according to the equation ha=h + a-.what are the concentrations of all the species in a solution containing 0.1 mole of ha per cubic decimetre?
1
Expert's answer
2019-09-20T07:36:34-0400

A weak acid dissociates according to the equation:

HA = H+ + A-

According to the acid dissociation constant (Ka) concept:

Ka = [H+] × [A-]/[HA],

where [H+], [A-], [HA] - molar concentrations of H+, A- and HA at equilibrium.

From here:

[HA] × Ka = [H+]×[A-]

As [H+] = [A-]:

[A-] = [H+] = ([HA] × Ka)1/2


[HA] = n/V,

where n - number of mokes (0.1 mole), V - volume (1 dm3 = 1L).

As a result, HA concentration equals:

[HA] = 0.1 mole / 1L = 0.1 M

Concentrations of [A-] and [H+] equal:

[A-] = [H+] = ([HA] × Ka)1/2 = (0.1 M × 10-9 )1/2 = 10-5 M


Answer: [HA] = 0.1 M, [H+] = 10-5 M, [A-] = 10-5 M.